Your car loan requires payments of $300 a month for the first year and payments of $500 per month the second year. The annual interest rate is 6%, and payments begin in one month. What is the present value of this two-year loan?
Answer #1

Calculating Present Value of first year payment,

Using TVM calculation,

PV = [FV = 0, PMT = 300, N = 12, I = 0.06/12]

PV = $3,485.68

Calculating Present Value(end of Year 1) of second year payment,

Using TVM Calculation,

PV = [FV = 0, PMT = 500, N = 12, I = 0.06/12]

PV = $5,809.47

Loan Amount = 3,485.68 + 5,809.47/(1.005)12

Loan Amount = $8,957.65